A shipment of 7 television sets contains 2 defective sets A hotel makes a random purchase of 3 of the sets: a) If x is the number of defective sets purchased by the hotel, find the probability distribution of x Express the results graphically as a probability histogram. b) Find the cumulative distribution function of the random variable X representing the number of defectives. Then using flx), find (a) p(X = 1); (b) p(O < X s 2).

1
Expert's answer
2021-11-13T02:10:01-0500

N=7

m=2

n=3

Let X be number of defective sets bought by the hotel

"P(x,N,n,m)=\frac{\begin{pmatrix}
 M \\
 x
\end{pmatrix}\begin{pmatrix}
 N-m \\
 n-x
\end{pmatrix}}{\begin{pmatrix}
 N \\
 n
\end{pmatrix}}"

"P(x,N,n,m)=\frac{\begin{pmatrix}
 2 \\
 0
\end{pmatrix}\begin{pmatrix}
 7-2 \\
 3-0
\end{pmatrix}}{\begin{pmatrix}
 7 \\
 3
\end{pmatrix}}"

"=\frac{\begin{pmatrix}
 2 \\
 0
\end{pmatrix}\begin{pmatrix}
 5 \\
 3
\end{pmatrix}}{\begin{pmatrix}
 7 \\
 3
\end{pmatrix}}"

"=\frac{(1)(10)}{(35)}"

=0.2857